OpenStudy (anonymous):

find the intervals for which f is increasing or decreasing and locate all relative extrema on (0,2pi). f(x)= cos x / 1 + sin x

OpenStudy (anonymous):

@satellite73

OpenStudy (science0229):

First thing to do: differentiate it.

OpenStudy (science0229):

I got -1/(sinx + 1)

OpenStudy (science0229):

Now, if the graph is increasing, the slope increases. If the graph decreases, the slope decreses. For any extrema, the slope is zero.

OpenStudy (science0229):

Solve the following 3 equations to get each answers. \[\frac{ -1 }{ \sin x + 1 }=0\] \[\frac{ -1 }{ \sin x + 1 } > 0\] \[\frac{ -1 }{ \sin x + 1 }<0\] where x is in the interval (0,2pi)

OpenStudy (science0229):

First equation for relative extrema Second equation for the interval the the graph increases Third equation for the interval that the graph decreases

OpenStudy (anonymous):

THANK YOU

OpenStudy (science0229):

Welcome :)

OpenStudy (anonymous):

@science0229

OpenStudy (anonymous):

I got another problem almost the same but instead its f(x)= cos x / 1 + sin ^ 2x

OpenStudy (anonymous):

could you help me?

OpenStudy (anonymous):

I will post a new question :)

OpenStudy (science0229):

Same process, with longer time. Differentiate it, first.